OpenStudy (sissyedgar):

How many solutions does the equation have? 4x + 2(x - 3) = 8x + 12 A. one solution B. infinite solutions C. no solution

OpenStudy (sissyedgar):

@imqwerty @Shalante @paki

OpenStudy (sissyedgar):

@sammixboo

OpenStudy (sissyedgar):

@Luigi0210 @mathlover2014 @kl0723

OpenStudy (sissyedgar):

@Abhisar @ganeshie8

OpenStudy (anonymous):

2(x-3)=2(x)-(2 x 3)=2x-6 4x+2x-6=8x+12 Combine like terms 6x-6=8x+12 Subtract 6x to both sides -6=2x+12 Subtract 12 from both sides -18=2x x=-9 Only one solution.

OpenStudy (sissyedgar):

YAY!!!!! ty!!!!!!!!